# Localized Deviations from the CO–PAH Relation in PHANGS–JWST Galaxies: Faint PAH Emission or Elevated CO Emissivity?

### v1

@@ Abstract
Polycyclic aromatic hydrocarbon (PAH) emission is widely used to trace the distribution of molecular gas in the interstellar medium (ISM), exhibiting a tight correlation with CO(2–1) emission across nearby galaxies. Using PHANGS-JWST and PHANGS-ALMA data, we identify localized regions where this correlation fails, with CO flux exceeding that predicted from 7.7m PAH emission by more than an order of magnitude. These outlier regions are found in 20 out of 70 galaxies and are mostly located in galaxy centers and bars, without signs of massive star formation. We explore two scenarios to explain the elevated CO-to-PAH ratios, which can either be due to suppressed PAH emission or enhanced CO emissivity. We examine PAH emission in other bands (3.3m and 11.3m) and the dust continuum dominated bands (10m and 21m), finding consistently high CO-to-PAH (or CO-to-dust continuum) emission ratios, suggesting that 7.7 m PAH emission is not particularly suppressed. In some outlier regions, PAH size distributions and spectral energy distribution of the illuminating radiation differ slightly compared to nearby control regions with normal CO-to-PAH ratios, though without a consistent trend. We find that the outlier regions generally show higher CO velocity dispersions (). This increase in may lower CO optical depth and raise its emissivity for a given gas mass. Our results favor a scenario where shear along the bar lanes and shocks at the tips of the bar elevate CO emissivity, leading to the breakdown of the CO–PAH correlation.
@@
1 Introduction
Emission from polycyclic aromatic hydrocarbons (PAHs) captured with JWST reveals a complex and widespread network of interstellar medium (ISM) structures throughout galaxy disks (J. C. Lee et al., 2023; D. A. Thilker et al., 2023; D. B. Fisher et al., 2024; T. G. Williams et al., 2024). In the diffuse Milky Way ISM, the fraction of cold neutral gas with respect to the total gas strongly correlates with PAH abundance, suggesting that PAHs are destroyed in warm gas while PAH growth and survival is facilitated in well-shielded, moderately dense gas (F. Galliano et al., 2018; B. S. Hensley & B. T. Draine, 2023a). In molecular gas-dominated galaxies, studies have shown a tight correspondence of PAH emission with CO emission as well as with H at scales ranging from a few kpc down to 100 pc (M. W. Regan et al., 2006; R. Chown et al., 2021; Y. Gao et al., 2022; A. K. Leroy et al., 2023; C. M. Whitcomb et al., 2023; R. Chown et al., 2025; H. Shim et al., 2025).
The tight correlation observed between CO and PAH emission in molecular gas-dominated galaxies can be understood within the standard framework of dust in galaxies, in which dust and PAH grains are well mixed within the gaseous phase of the ISM (B. T. Draine & A. Li, 2007; F. Galliano et al., 2018; B. S. Hensley & B. T. Draine, 2023a). PAHs are stochastically excited by ultraviolet and optical photons, allowing them to emit efficiently in the mid-infrared. As a result, the distribution of mid-IR emission is expected, to first order, to closely resemble the distribution of the cold ISM (A. K. Leroy et al., 2023; R. Chown et al., 2025). Indeed, PAH emission has even been used to probe the ISM in Hi-dominated regions with gas surface densities below (K. M. Sandstrom et al., 2023a).
@@ 1 Introduction
Recently, R. Chown et al. (2025) have characterized the relationship between 12CO(; hereafter CO) and PAH emission across 70 galaxies from the Physics at High Angular Resolution in Nearby Galaxies Survey (PHANGS). By combining Atacama Large Millimeter/sub-millimeter Array (ALMA) CO maps and JWST near- and mid-infrared (near-IR and mid-IR) imaging, R. Chown et al. (2025) have found a relatively small scatter () in the observed CO-to-PAH ratios at 100 pc-scale, where CO shows a strong and approximately linear correlation with 7.7m PAH emission, one of the common PAH tracers accessible with JWST. Similarly linear trends were observed at 11.3m and 3.3m for a subsample of 19 galaxies from the PHANGS–JWST Cycle 1 survey (J. C. Lee et al., 2023).
@@
The growing evidence of PAH closely tracing CO has led to the suggestion that PAH emission can be used as a tracer of the cold ISM, not only for high–resolution studies of nearby galaxies but also out to high redshift (I. Cortzen et al., 2019; I. Shivaei & L. A. Boogaard, 2024; R. Chown et al., 2025; H. Shim et al., 2025). This prospect is particularly exciting because JWST is now routinely enabling high–resolution, high–sensitivity PAH imaging. Looking ahead, Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer satellite (SPHEREx; O. Doré et al., 2018) will deliver full–sky spectral PAH maps, opening the door to uniform cold ISM mapping across all nearby galaxies. However, for these applications to be reliable, we need to identify where the relation between PAH and cold gas breaks down, beyond classical H II regions, for example in AGN-dominated centers and zones of strong shear or shocks.
@@
In this Letter, we examine regions that deviates significantly from the CO–7.7m PAH relation as calibrated by R. Chown et al. (2025) across the same PHANGS galaxies. We then assess the local environment, galactic dynamics, and excitation conditions of these outliers to isolate the physical processes that drive these deviations from the global correlation.
2 Observational Data
@@
We focus on the same 70 galaxies from R. Chown et al. (2025) for which scaling relations between CO and 7.7m PAH emission were derived using high-resolution JWST imaging with the F770W filter and ALMA CO observations. These targets are from PHANGS–JWST Cycle 1 (GO 2107; PI: Lee; J. C. Lee et al., 2023; T. G. Williams et al., 2024) and Cycle 2 (GO 3707; PI: Leroy; R. Chown et al., 2025). Most of these galaxies lie on the star-forming main sequence, with stellar masses ranging from to 11.5 and at distances of Mpc.
2.1 JWST mid-IR imaging of PAH emission
@@
For 19 galaxies from the PHANGS-JWST Cycle 1 program, we use JWST NIRCam and MIRI imaging data observed with F335M, F770W, F1000W, F1130W, and F2100W filters. The Cycle 1 data can be obtained using this link111https://archive.stsci.edu/hlsp/phangs or Williams, Thomas et al. (2023). The remaining 51 galaxies are from the PHANGS-JWST Cycle 2 program described in R. Chown et al. (2025), with limited coverage in filters. We use two mid-IR bands of F770W and F2100W.
@@
The F335M, F770W and F1130W filters capture strong PAH emission features at 3.3, 7.7 and 11.3m. The F1000W (centered at 10m) is expected to be dominated by dust continuum (C. M. Whitcomb et al., 2023), though its morphology and intensity often track CO and H, similar to the PAH bands (A. K. Leroy et al., 2023). The F2100W at 21m captures dust continuum emission, including contributions from grains in thermal equilibrium and heated by the interstellar radiation field (B. T. Draine & A. Li, 2007; R. C. Kennicutt & N. J. Evans, 2012; A. K. Leroy et al., 2012; F. Belfiore et al., 2023a; H. Hassani et al., 2023; A. K. Leroy et al., 2023; D. Pathak et al., 2024). The resolution of F335M, F770W, F1000W, F1130W, F2100W imaging is 011, 027, 033, 038, and 067, respectively222Point spread functions for NIRCAM and MIRI, where the resolution of F2100W roughly translates into a physical scale of pc across the range of distances of our galaxy sample ( Mpc). Both of the Cycle 1 and Cycle 2 data have been reduced using the PHANGS-JWST pipeline (pjpipe333https://pjpipe.readthedocs.io/en/latest/). For comprehensive descriptions of the survey and the data reduction process, we refer readers to J. C. Lee et al. (2023), T. G. Williams et al. (2024), and R. Chown et al. (2025).
@@
The F335M and F770W bands include contributions from the stellar continuum. For F335M, we use continuum-subtracted maps from H. Koziol et al. (in prep.), where the stellar continuum is estimated using F300M and F360M, based on the method of K. M. Sandstrom et al. (2023b). As for F770W, we used stellar continuum-subtracted F770W maps adopted in R. Chown et al. (2025), obtained by subtracting the flux in F200W (Cycle 1) or F300M (Cycle 2) scaled by a factor determined using the prescription of J. Sutter et al. (2024).
2.2 ALMA 12CO( imaging
@@
We use 12CO() imaging from the PHANGS–ALMA survey, hereafter CO. Full descriptions of the sample, survey design, and image production process are provided in A. K. Leroy et al. (2021a) and A. K. Leroy et al. (2021b). The observations were conducted with the 12-m, 7-m, and total power arrays of ALMA. The CO maps have resolutions of 10-17, corresponding to physical scales of 25–150 pc for the galaxies in our sample. We use the first public release of the moment-0 maps, generated using an inclusive signal-masking scheme (“broad” masking; see A. K. Leroy et al., 2021b) designed to capture extended and faint emission and encompass all emission from each galaxy, at the native resolution. We also use maps of the CO effective line width () from the same public release, produced with the more restrictive “strict” mask, which includes only voxels highly likely to contain real emission.
2.3 ALMA 12CO( imaging
@@
2.4 MUSE ionized gas emission
@@
2.5 Convolution and homogenization to a common pixel grid
@@
3 Outlier detection
We perform outlier detection using CO and stellar continuum subtracted F770W maps that has been resampled on a gird with one CO-beam spacing to reduce inter-sample covariance. We then identify pixels whose CO emission is anomalously high relative to the 7.7 m PAH emission. For each galaxy, we predict CO emission from the stellar–continuum–subtracted F770W emission () using the relation:
where the coefficients , , and are individual galaxy based fits from Table 4 of R. Chown et al. (2025). We then calculate residuals
@@ 3 Outlier detection
where and are per beam measurement errors, are the fit–parameter uncertainties, and is the intrinsic scatter of the relation from R. Chown et al. (2025).
To remain consistent with R. Chown et al. (2025), we limit the analysis to beam samples with . To avoid noise contamination and isolate significant outliers, we additionally require in both and . We then identify beam samples with and group adjacent beams into a contiguous region and compute inverse–variance–weighted statistics to calculate average residual and its uncertainty ( and ). We define a contiguous region as having abnormally high CO-to-PAH ratio if it satisfies the conservative detection threshold of . This results in a detection of outlier regions in galaxies out of 70 galaxies analyzed in R. Chown et al. (2025). If we relax the threshold to , additional outlier regions that are relatively small, composed of two to three CO beams, emerge in NGC 1512, NGC 1672, NGC 2835, NGC 4457, and NGC 4548. Throughout the paper, we adopt a more conservative threshold to focus on significant and obvious outliers.
@@
In Figure 1 we show the measured ratio () for three galaxies, selected because their outlier regions are either spatially extended or located across diverse environments. The red contours in Figure 1 highlight regions where the observed CO flux exceeds the value expected from the 7.7m PAH emission by more than an order of magnitude. In Appendix A, we present the same figure for all 20 galaxies with such outliers, along with their properties, among the 70 galaxies included in our analysis. We note that most galaxies hosting outliers are classified as strongly or weakly barred (17 out of 20; R. J. Buta et al. 2015). However, outliers are not ubiquitous among barred systems, with only 35% of barred galaxies exhibiting outliers.
These abnormal regions represent significant outliers from the tight CO–7.7m PAH relation reported by R. Chown et al. (2025). They constitute 0.03-7% of the total CO-emitting area with an average of 1% (see Appendix A) and are spatially localized along bar lanes and centers without signs of massive star formation detected in H. NGC 1566 is the only galaxy where we find such regions at the edges of spiral arms. The preferential location of these regions within specific dynamical structures suggests that they are not noise-driven statistical outliers. Instead, the elevated CO-to-PAH ratios have a physical origin linked to local gas dynamics and/or radiation environments.
In order to investigate the origin of these high CO-to-PAH ratio regions, we define a “null test” control region with three requirements: 1) it lies near the outlier region within the same dynamical structure; 2) it contains only pixels with normal CO-to-PAH ratios (); and 3) it shows no obvious signs of recent star formation, so as to remain consistent with the outlier regions. Recent star formation is traced by extinction-corrected MUSE H for Cycle 1 galaxies and by narrowband H imaging for Cycle 2 galaxies. In addition to these, whenever possible, we also required the null test control region to have a CO flux comparable to that of the outlier region. The locations of these null test regions are shown in Appendix A.
@@
4 Origin of outliers
4.1 Suppressed PAH Emission Scenario

### v2

@@ Abstract
Polycyclic aromatic hydrocarbon (PAH) emission is widely used to trace the distribution of molecular gas in the interstellar medium, exhibiting a tight correlation with CO(2–1) emission across nearby galaxies. Using PHANGS-JWST and PHANGS-ALMA data, we identify localized regions where this correlation fails, with CO flux exceeding that predicted from 7.7m PAH emission by more than an order of magnitude. These outlier regions are found in 20 out of 70 galaxies and are located in galaxy centers and bars, without signs of massive star formation. We explore two scenarios to explain the elevated CO-to-PAH ratios, which can either be due to suppressed PAH emission or enhanced CO emissivity. We examine PAH emission in other bands (3.3m and 11.3m) and the dust continuum dominated bands (10m and 21m), finding consistently high CO-to-PAH (or CO-to-dust continuum) emission ratios, suggesting that 7.7 m PAH emission is not particularly suppressed. In some outlier regions, PAH sizes and spectral energy distribution of the radiation differ slightly from nearby control regions with normal CO-to-PAH ratios, though without a consistent trend. We find that the outlier regions show higher CO velocity dispersions (). This increase in lowers CO optical depth and raises its emissivity for a given gas mass. Our results favor a scenario where shear along the bar lanes and shocks at the bar ends elevate CO emissivity, leading to the breakdown of the CO–PAH correlation. Future JWST spectroscopy and deep ALMA observations of CO isotopologues will provide critical tests of this scenario.
@@
I Introduction
Emission from polycyclic aromatic hydrocarbons (PAHs) captured with JWST reveals a complex and widespread network of interstellar medium (ISM) structures throughout galaxy disks (J. C. Lee et al., 2023; D. A. Thilker et al., 2023; D. B. Fisher et al., 2024; T. G. Williams et al., 2024). In the diffuse Milky Way ISM, the fraction of cold neutral gas with respect to the total gas strongly correlates with PAH abundance, suggesting that PAHs are destroyed in warm gas while PAH growth and survival is facilitated in well-shielded, moderately dense gas (F. Galliano et al., 2018; B. S. Hensley & B. T. Draine, 2023a). In molecular gas-dominated galaxies, studies have shown a tight correspondence of PAH emission with CO emission as well as with H at scales ranging from a few kpc down to 100 pc (M. W. Regan et al., 2006; R. Chown et al., 2021; Y. Gao et al., 2022; A. K. Leroy et al., 2023; C. M. Whitcomb et al., 2023; R. Chown et al., 2025a; H. Shim et al., 2025).
The tight correlation observed between CO and PAH emission in molecular gas-dominated galaxies can be understood within the standard framework of dust in galaxies, in which dust and PAH grains are well mixed within the gaseous phase of the ISM (B. T. Draine & A. Li, 2007; F. Galliano et al., 2018; B. S. Hensley & B. T. Draine, 2023a). PAHs are stochastically excited by ultraviolet and optical photons, allowing them to emit efficiently in the mid-infrared. As a result, the distribution of mid-IR emission is expected, to first order, to closely resemble the distribution of the cold ISM (A. K. Leroy et al., 2023; R. Chown et al., 2025a). Indeed, PAH emission has even been used to probe the ISM in Hi-dominated regions with gas surface densities below (K. M. Sandstrom et al., 2023a).
@@ I Introduction
Recently, R. Chown et al. (2025a) have characterized the relationship between 12CO(; hereafter CO) and PAH emission across 70 galaxies from the Physics at High Angular Resolution in Nearby Galaxies Survey (PHANGS). By combining Atacama Large Millimeter/sub-millimeter Array (ALMA) CO maps and JWST near- and mid-infrared (near-IR and mid-IR) imaging, R. Chown et al. (2025a) have found a relatively small scatter () in the observed CO-to-PAH ratios at 100 pc-scale, where CO shows a strong and approximately linear correlation with 7.7m PAH emission, one of the common PAH tracers accessible with JWST. Similarly linear trends were observed at 11.3m and 3.3m for a subsample of 19 galaxies from the PHANGS–JWST Cycle 1 survey (J. C. Lee et al., 2023).
@@
The growing evidence of PAH closely tracing CO has led to the suggestion that PAH emission can be used as a tracer of the cold ISM, not only for high–resolution studies of nearby galaxies but also out to high redshift (I. Cortzen et al., 2019; I. Shivaei & L. A. Boogaard, 2024; R. Chown et al., 2025a; H. Shim et al., 2025). This prospect is particularly exciting because JWST is now routinely enabling high–resolution, high–sensitivity PAH imaging. Looking ahead, Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer satellite (SPHEREx; O. Doré et al., 2018) will deliver full–sky spectral PAH maps, opening the door to uniform cold ISM mapping across all nearby galaxies. However, for these applications to be reliable, we need to identify where the relation between PAH and cold gas breaks down, beyond classical Hii regions, for example in AGN-dominated centers and zones of strong shear or shocks.
@@
In this Letter, we examine regions that deviates significantly from the CO–7.7m PAH relation as calibrated by R. Chown et al. (2025a) across the same PHANGS galaxies. We then assess the local environment, galactic dynamics, and excitation conditions of these outliers to isolate the physical processes that drive these deviations from the global correlation.
II Observational Data
@@
We focus on the same 70 galaxies from R. Chown et al. (2025a) for which scaling relations between CO and 7.7m PAH emission were derived using high-resolution JWST imaging with the F770W filter and ALMA CO observations. These targets are from PHANGS–JWST Cycle 1 (GO 2107; PI: Lee; J. C. Lee et al., 2023; T. G. Williams et al., 2024) and Cycle 2 (GO 3707; PI: Leroy; R. Chown et al., 2025a). Most of these galaxies lie on the star-forming main sequence, with stellar masses ranging from to 11.5 and at distances of Mpc.
II.1 JWST mid-IR imaging of PAH emission
@@
For 19 galaxies from the PHANGS-JWST Cycle 1 program, we use JWST NIRCam and MIRI imaging data observed with F335M, F770W, F1000W, F1130W, and F2100W filters. The Cycle 1 data can be obtained using this link111https://archive.stsci.edu/hlsp/phangs or Williams, Thomas et al. (2023). The remaining 51 galaxies are from the PHANGS-JWST Cycle 2 program described in R. Chown et al. (2025a), with limited coverage in filters. We use two mid-IR bands of F770W and F2100W.
@@
The F335M, F770W and F1130W filters capture strong PAH emission features at 3.3, 7.7 and 11.3m. The F1000W (centered at 10m) is expected to be dominated by dust continuum (C. M. Whitcomb et al., 2023), though its morphology and intensity often track CO and H, similar to the PAH bands (A. K. Leroy et al., 2023). The F2100W at 21m captures dust continuum emission, including contributions from grains in thermal equilibrium and heated by the interstellar radiation field (B. T. Draine & A. Li, 2007; R. C. Kennicutt & N. J. Evans, 2012; A. K. Leroy et al., 2012; F. Belfiore et al., 2023a; H. Hassani et al., 2023; A. K. Leroy et al., 2023; D. Pathak et al., 2024). The resolution of F335M, F770W, F1000W, F1130W, F2100W imaging is 011, 027, 033, 038, and 067, respectively222Point spread functions for NIRCAM and MIRI, where the resolution of F2100W roughly translates into a physical scale of pc across the range of distances of our galaxy sample ( Mpc). Both of the Cycle 1 and Cycle 2 data have been reduced using the PHANGS-JWST pipeline (pjpipe333https://pjpipe.readthedocs.io/en/latest/). For comprehensive descriptions of the survey and the data reduction process, we refer readers to J. C. Lee et al. (2023), T. G. Williams et al. (2024), and R. Chown et al. (2025a).
@@
The F335M and F770W bands include contributions from the stellar continuum. For F335M, we use continuum-subtracted maps from H. Koziol et al. (in prep.), where the stellar continuum is estimated using F300M and F360M, based on the method of K. M. Sandstrom et al. (2023b). As for F770W, we used stellar continuum-subtracted F770W maps adopted in R. Chown et al. (2025a), obtained by subtracting the flux in F200W (Cycle 1) or F300M (Cycle 2) scaled by a factor determined using the prescription of J. Sutter et al. (2024).
II.2 ALMA 12CO( imaging
@@
We use 12CO() imaging from the PHANGS–ALMA survey, hereafter CO. Full descriptions of the sample, survey design, and image production process are provided in A. K. Leroy et al. (2021a) and A. K. Leroy et al. (2021b). The observations were conducted with the 12-m, 7-m, and total power arrays of ALMA. The CO maps have resolutions of 10-17, corresponding to physical scales of 25–150 pc for the galaxies in our sample. We use the first public release of the moment-0 maps, generated using an inclusive signal-masking scheme (“broad” masking; see A. K. Leroy et al., 2021b), designed to capture extended and faint emission and encompass all emission from each galaxy at the native resolution. We also use maps of the CO effective line width () from the same public release, produced with the more restrictive “strict” mask, which includes only voxels highly likely to contain real emission.
II.3 ALMA 12CO( imaging
@@
II.4 MUSE ionized gas emission
@@
II.5 Convolution and homogenization to a common pixel grid
@@
III Outlier detection
We perform outlier detection using CO and stellar continuum subtracted F770W maps that has been resampled on a gird with one CO-beam spacing to reduce inter-sample covariance. We then identify pixels whose CO emission is anomalously high relative to that expected from the 7.7 m PAH emission. For each galaxy, we predict CO emission from the stellar–continuum–subtracted F770W emission () using the relation:
where the coefficients , , and are individual galaxy based fits from Table 4 of R. Chown et al. (2025a). We then calculate residuals
@@ III Outlier detection
where and are per beam measurement errors, are the fit–parameter uncertainties, and is the intrinsic scatter of the relation from R. Chown et al. (2025a).
To remain consistent with R. Chown et al. (2025a), we limit the analysis to beam samples with . To avoid noise contamination and isolate significant outliers, we additionally require in both and . We then identify beam samples with and group adjacent beams into a contiguous region and compute inverse–variance–weighted statistics to calculate average residual and its uncertainty ( and ). We define a contiguous region as having abnormally high CO-to-PAH ratio if it satisfies the conservative detection threshold of . This results in a detection of outlier regions in 20 galaxies out of 70 galaxies analyzed in R. Chown et al. (2025a). If we relax the threshold to , additional outlier regions that are relatively small, composed of two to three CO beams, emerge in NGC 1512, NGC 1672, NGC 2835, NGC 4457, and NGC 4548. Throughout the paper, we adopt a more conservative threshold to focus on significant and obvious outliers.
As an additional robustness check, we varied the flux ratio criterion applied to beam samples from to . Lowering the cut to 8 identifies additional outlier regions in NGC 1512, NGC 1672, NGC 1792, NGC 4457, NGC 4548, and NGC 7496. On the other hand, adopting a stricter cut of 12 removes smaller outliers and causes a few galaxies that previously hosted outliers to no longer meet the criterion (NGC 3351, NGC 4321, NGC 4731, and NGC 4941). Nonetheless, the key qualitative trends and our main conclusions remain unchanged.
@@
In Figure 1, we show the measured ratio () for three galaxies, selected because their outlier regions are either spatially extended or located across diverse environments. The red contours in Figure 1 highlight regions where the observed CO flux exceeds the value expected from the 7.7m PAH emission by more than an order of magnitude. In Appendix A, we present the same figure for all 20 galaxies with such outliers, along with their properties in Table 1, among the 70 galaxies included in our analysis. We note that most galaxies hosting outliers are classified as strongly or weakly barred (17 out of 20; R. J. Buta et al. 2015). However, outliers are not ubiquitous among barred systems, with only 35% of barred galaxies exhibiting outliers.
These outlier regions represent significant deviations from the tight CO–7.7m PAH relation reported by R. Chown et al. (2025a). They constitute 0.03-7% of the total CO-emitting area with an average of 1% (see Appendix A) and are spatially localized along bar lanes and centers without signs of massive star formation detected in H. NGC 1566 is the only galaxy where we find such regions at the edges of spiral arms. The preferential location of these regions within specific dynamical structures suggests that they are not noise-driven statistical outliers. Instead, the elevated CO-to-PAH ratios have a physical origin linked to local gas dynamics and/or radiation environments.
In order to investigate the origin of these high CO-to-PAH ratio regions, we define a “null-test” control region with three requirements: 1) it lies near the outlier region within the same dynamical structure; 2) it contains only pixels with normal CO-to-PAH ratios (); and 3) it shows no obvious signs of recent star formation, so as to remain consistent with the outlier regions. Recent star formation is traced by extinction-corrected MUSE H for Cycle 1 galaxies and by narrowband H imaging for Cycle 2 galaxies. In addition to these, whenever possible, we also required the null-test control region to have a CO flux comparable to that of the outlier region. In practice, we select null-test regions near the outlier regions within the same dynamical structure, which closely matches the local environment and the galactocentric radius. This also allows us to mitigate the effect of metallicity gradients in our comparison between null-test and outlier region555Even if there is some metallicity differences between the null-test and outlier regions, both CO and PAH emission are expected to increase at higher metallicity due to a more efficient shielding and a higher PAH fraction (e.g., C. W. Engelbracht et al., 2005; R. Chown et al., 2025b), making metallicity variations an unlikely driver of the most extreme CO/PAH outliers.. In some cases we do not choose the immediately neighboring area in order to avoid star-forming sites, but we still aim to select regions with comparable CO flux. The null-test regions therefore remain at similar galactocentric radii, show no obvious signs of recent star formation, and have sufficient CO signal for a meaningful comparison. The locations of these null test regions are shown in Appendix A.
@@
IV Origin of outliers
IV.1 Suppressed PAH Emission Scenario